Thread View: datac
1 messages
1 total messages
Started by utzoohenry
Thu, 04 Feb 1982 15:47
hardware status
Author: utzoohenry
Date: Thu, 04 Feb 1982 15:47
Date: Thu, 04 Feb 1982 15:47
6 lines
563 bytes
563 bytes
Wern has switched the floppies from the 11/05 for the ones he's got on the 11/34; this "fixes" the floppy drive that has been dead for the last few days. He has also found and fixed the cable problem that inactivated the Pomz and Dunham crates, and more-or-less cured the power-supply problem that had Star Ring down. Star Ring is not yet back on Unix, this will probably happen tonight.
Thread Navigation
This is a paginated view of messages in the thread with full content displayed inline.
Messages are displayed in chronological order, with the original post highlighted in green.
Use pagination controls to navigate through all messages in large threads.
Back to All Threads